OpenStudy (anonymous):

what is the area of a square inscribed in a circle with radius 6?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

first draw the circle inside the square . Now we all know that the relationship between the diameter and R is : R = D/2 , so D = 2R If you find the diameter of the circle, you'll notice that it is equal to the side of the sqaure D = S , so after finding the diameter of the circle , substitute it in the general formula of the square's area which is : \[A = S^2\] ^_^ that's it

OpenStudy (anonymous):

square*